Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find P80.


1
Expert's answer
2022-02-08T16:50:10-0500

Assume the data is:

0.1

0.2

0.3

0.5

0.6

0.6

0.7

0.7

0.8

0.8

0.8

0.8

0.8

0.9

0.9

1.3

1.4

1.5

1.6

1.6

1.6

2.1

2.1

2.4

2.4

2.5

3.4

3.6

4.1

4.2

4.5

4.6

5.7

6.7

7.7

8.1

8.5

12.6

12.6

12.8

12.8

13.3

13.3

13.7

14.2

14.3

15.1

15.5

15.6

28.6


Solution:

The data are in ascending order.

Compute the position of the pth percentile (index "i" ):



"i=\dfrac{p}{100}\times n=\dfrac{80}{100}\times50=40"

The index "i" is an integer ⇒ the 80th percentile is the average of the values in the 39th and 40th positions (12.8 and 12.8 respectively)

Answer: the 80th percentile is (12.8 + 12.8) / 2 = 12.8


P80=1.6 Mbps